How to find ten consecutive even integers that sum to 70 and have a product of 0. Determine the product of the largest and smallest even integers in this set. Explain how did I get the answer and show work.
Answer by ankor@dixie-net.com(22740)

How to find ten consecutive even integers that sum to 70 and have a product of 0. Determine the product of the largest and smallest even integers in this set. Explain how did I get the answer and show work.
:
The sum of 10 consecutive even integers, x, (x+2), (x+4), (x+6) etc
10x + 90 = 70
10x = 70 - 90
10x = -20
x = -2
:
-2*0*2*4*6*8*10*12*14*16 = 0, with a zero in there the product will always be 0
:
Product of the smallest and largest: -2 * 16 = -32